Assume that you have a nozzle, with air entering at 300 kPa and 67 C with a velocity of 60 m/s. The air exits the nozzle at 90 kPa and 330m/s. Do NOT make an adiabatic assumption. Instead, assume that the heat losses from the nozzle to the surroundings are 3.2 kJ/kg. The surroundings are at 20 C. Assume that air is an ideal gas with variable specific heats

.a) Find the exit temperature.

b) Find the entropy change for the air (kJ/kg K)

c) Find the total entropy change(kJ/kg K). (Consider both the air in the nozzle and the surroundings.)
Answer #1

02 P = sookpad T-67°c P2 = go (kPa) V2 = $50 (18) = 340K V = 60lm 18) 2 By Applying 21-22 assumed to an SFEE between Of ® e hI sa-si = -0.16817014 + 0.3u55uolg * 52-51 (pen (1) - Ren (1) fligg) & S2-51 0.287x en/287 611gy! 1.005 -0.287 x 90 300 340 0